# Toxicity Detection is NOT all you Need: Measuring the Gaps to Supporting Volunteer Content Moderators through a User-Centric Method

### v2

@@ 1 Introduction
In this paper, we aim to assess to what extent current natural language processing (NLP) models can serve the wide spectrum of moderation rules that exist. First, to understand the functions previous automated content moderation models have focused on, we conduct a model review on Hugging Face (HF) with rules from three subreddits as exemplars. This allows us to gauge the progress of past model developments in covering various moderation rules. We use model review as opposed to the more common literature review to gain a technical understanding of the existing models’ functions. In addition to examining models that are built to handle specific tasks, we also assess so-called “general-purposed” large language models’ (LLMs’) capability in covering various moderation rules. We evaluate GPT-4 and Llama-2 on a new evaluation dataset that we collected, consisting of moderation decisions from r/AskHistorians and covering a wide range of rules. Finally, using the models’ performance on this new dataset as an empirical grounding, we conduct a survey study with r/AskHistorians moderators ($N=11$) to evaluate the models’ performance and gain insights on performance requirements for a useful model for different rules.
We find a substantial gap in both existing NLP models and LLMs’ performance when to cover the diverse set of moderation rules that subreddits employ. Our analysis shows the majority of moderation rules from the three subreddits (approximately $80\%$) are unrelated to toxicity detection, with nearly $70\%$ lacking an huggingface model designed for their resolution. While one might hope that general-purpose LLMs could fill this gap, our experiments with GPT-4 and Llama-2 show that LLMs fall short in their ability to detect violations of many rules. Specifically, both GPT-4 and Llama-2 exhibited moderate to low precision and/or recall ($<70\%$) for half of the rules from r/AskHistorians. Findings from our survey study also indicate that neither LLM has good enough performance to be useful for $6$ of $23$ r/AskHistorians rules ($26\%$). Finally, our show survey study shows that moderators are excited about an assistant model but have complex needs in terms of model precision and recall for different kinds of rules. Our study highlights the necessity of, and provides a guide for, future content moderation work to expand its focus beyond toxicity detection, encompassing a broader spectrum of moderation rules to meet the needs of moderators seeking automation assistance.
@@ 2 Background and Related Work
Recently, Kumar et al. (2023) tested LLMs’ performance on rule-based moderation with rules from $95$ subreddits. They evaluated GPT models (3, 3.5, 4) on a dataset consisting of removed Reddit posts as violating and not violating content. They conclude that the GPT models are effective in conducting moderation on subreddits like r/Movies but struggle with other subreddits such as r/askscience and r/AskHistorians Our study builds on previous results by studying gaps in both purpose-built moderation models and LLMs. Rather than focusing on the broad binary questions of removing or keeping posts, we focus on coverage for specific moderation rules. In addition, we include the perspectives of human moderators to understand how existing models address the needs of real-world moderation.
@@ 3.1 Method
We then search on HF by matching keywords with model names or model cards to find the most relevant model to each moderation rule. We traverse the top 20 search results to find the most relevant model, which we call the matching model. If more than one model is relevant, we select the model with the highest number of likes on HF. If nothing relevant is found, we declare there is no matching model for this moderation rule. We also skip any model that does not contain a model card (i.e., no model description). See Tables 3, 4, and 5 in the appendix for the keywords we used and the matching models for each rule.
@@ 3.2 Model Review Results
Among the $37$ non-toxicity-related rules, only $6$ ($20\%$) rules have matching HF models that can be applied directly. $13$ ($43\%$) of the rules require some model modifications in order for the matching model to be adapted for the rule, whereas $11$ ($37\%$) rules have no matching model at all. These rules include, for example, low-effort posts and proselytizing from r/Atheism, and no homework and no ‘‘soapboxing’’ or loaded questions from r/AskHistorians.
@@
To see detailed annotations for each question, please refer to Tables 3, 4, and 5 in appendix.
@@ 4.1 Method
The evaluation dataset was created and annotated by a volunteer moderator for r/AskHistorians who has expertise in both moderation and qualitative analysis. For each rule, 1111111113131313 violating posts or comments were selected, save for posts violating the Illegal artifacts rule, which, due to the rareness of violations, was only represented $3$ times in the dataset. Most content was selected from r/AskHistory, a Reddit community that is thematically similar to r/AskHistorians but with different moderation rules. Some additional content was taken directly from r/AskHistorians. We exclude all borderline cases and only select posts and comments that clearly violate the rules. In total, the dataset includes $101$ questions and $134$ comments. [^4] In r/AskHistorians, posts are questions users have asked and comments are responses to questions. Comments included in the dataset are not necessarily responses to the questions included in the dataset. Because content often violates more than one rule, questions and comments included in the dataset often reflect multiple violations; however, annotations represent the most relevant or severe violation. For example, a comment containing only a joke would be annotated as violating the rule prohibiting joke responses, even though it would also violate rules on comprehensiveness or depth. The dataset also included $11$ questions and $10$ comments that do not violate any r/AskHistorians rules. We use these data as our evaluation dataset. [^5] The dataset will be published.
@@
Finally, following the suggestions from Shieh (2023), we set the model’s role as a moderator assistant and put the main question at the beginning of the prompt. In the few-shot setting, we also repeat the main question at the end of the prompt. See Appendix A for the exact prompts we used.
@@ 4.2 Moderator Evaluation
To understand volunteer moderators’ perspectives on the relative performance of LLMs, we also conducted an evaluation survey with active moderators from r/AskHistorians. [^6] Approved by our institutional IRB. The survey aimed to capture 1) whether moderators would use the LLM to help with their moderation work on each rule, given the LLM’s performance (precision and recall) identifying violations for this rule, and 2) for each rule, what kind of model performance is needed to effectively assist moderators.
@@ 4.3 Model Performance – Results
Our survey results indicate that different moderators have different perceptions of what constitutes an adequate model: for the evaluation questions, participants exhibited only fair agreement (Fleiss’ Kappa $0.34$). Mirroring policy decision-making standards in the subreddit, we use majority vote to summarize participants’ evaluation of the models’ performance. The results are shown in Figure 2 and Figure 3. In most cases, moderators would consider models with $>70\%$ precision and recall to be useful. Among the $23$ rules, both models are considered useful for three question rules and three comment rules. For some rules, such as NSFW/Sensitive content, No plagiarism, No political agendas or moralising, and Depth, moderators can accept low recall; for some, such as Do not just post links or quotations, moderators can accept low precision. Importantly, there are six rules ($26\%$ of all rules) for which neither model is considered useful. This emphasizes that, while current LLMs provide possibilities for supporting content moderation work, further improvement is needed to support useful moderation tools.
@@ 5 Discussion
Furthermore, we observed varying requirements among users and rules. Even for the same rule, different moderators expressed distinct preferences for high precision or recall. This suggests that in situations where a trade-off between precision and recall is necessary, moderators may lean towards a model offering more user control, thus allowing for an adjustable trade-off. This way, they can customize the model to align with their preferences. In addition, for complex rules (e.g. Plagiarism and Digression) or subjective rules (e.g. Basic facts and Scope), moderators prefer model explanations over mere flagging of violations.
@@ 7 Limitations
Appendix A Prompt for Testing LLMs
A.1 Prompt for Testing Llama-2 with the Few-shot Setting
@@
A.2 Prompt for Testing GPT-44 with the Zero-shot Setting
@@
Appendix B LLMs Performances under Zero-shot and Few-shot Settings
@@
Appendix C User Survey
@@
Advertising & Commercialization
No Spam & Self-promotion - Same Source Posts
@@
Content/Behavior
Homework; No "what if" questions Basic facts Depth; No speculation; Jokes and humor
Proselytizing
No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs; No Extraneous Comic Book Movie Submission"
Copyright
No plagiarism
Doxxing/Personal Info
Privacy; No personal anecdotes
Personal Attacks or Flaming
Format
No partial answers or "placeholders"; Do not just post links or quotations
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title
Harassment
Civility
Harassment or Bigotry
No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
Hate Speech
Civility
No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
@@
Links & Outside Content
Sources
Prohibited Popular Websites
Low-Quality Content
Scope; Clutter
Low-Effort Posts
NSFW
NSFW/Sensitive content
Off-topic
Digression
Off Topic
No Extraneous Comic Book Movie Submission
Personal Army
Brigading
No Subreddit Brigading
Personality
Civility
Personal Attacks or Flaming
No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
Politics
No political agendas or moralising; No "Soapboxing" or loaded questions
@@
Reposting
No Repost or Discussion Threads of New Releases
@@
Spam
Spam
No Spam & Self-promotion - Same Source Posts
Spoilers
No Repost or Discussion Threads of New Releases; No Spoilers
Trolling
No Trolling
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title
Voting
"Poll-type" question
Others
Current event; Illegal artifacts
Don’t complain about the use of AAVE or slang
No Encouraging Piracy
@@
All users are expected to behave with courtesy and politeness at all times. We will not tolerate racism, sexism, or any other forms of bigotry. This includes Holocaust denialism. Nor will we accept personal insults of any kind, and do not allow minor nitpicking of grammar or spelling.
Scope
@@
Submissions must be about a question about the human past, a META post about the state of the subreddit, or an AMA ("Ask Me Anything") with a historical expert or panel of experts.
No Current events
@@
To discourage off-topic discussions of current events, questions, answers, and all other comments must be confined to events that happened 20 years ago or more, inclusively (e.g., 2003 and older).
Homework
@@
Our users aren’t here to do your homework for you, but they might be willing to help. Don’t just give us your essay/assignment topic and ask us for ideas.
NSFW/Sensitive content
@@
Questions with NSFW titles will be deleted, and we will ask you to repost it with a different title.
No "Poll"-type questions
@@
"Poll"-type questions aren’t appropriate here: "Who was the most influential person in history?" or "Who was the worst general in your period?" or "Who are your Top 10 favorite people in history?" If your question includes the words "most" or "least", or "best" or "worst" (or can be reworded to include these words), it’s probably a "poll"-type question.
No "Soapboxing" or Loaded Questions
@@
All questions must allow a back-and-forth dialogue based on the desire to gain further information and not be predicated on a false and loaded premise in order to push an agenda. Example: Good Question: "People say that Nixon is the worst President of all time. Why is this so?" Bad Question: "Nixon was the worst President of all time. Why isn’t Obama considered the worst?" The bad question is a fishing expedition to try to start a debate about Obama’s presidency. Most of these questions will break our 20-year rule, or try to set up a debate about an issue using a long wall of text in the main post.
No "What If" Questions
@@
Questions should be about what did happen, not what could have happened.
Privacy
@@
We do not allow questions that pose possible privacy issues for living or recently deceased persons who are not in the public eye. The cut-off for "recent" is 100 years.
Illegal artifacts
@@
It is our policy to disallow posts asking for further information on artifacts where there is a likelihood that the acquisition or possession of the item might be illegal, unethical, and/or run contrary to sound, historical practices.
Basic Facts
@@
Questions looking for specific, basic facts - for the purpose of this rule, seeking a name, a date or time, a number, a location, the origin of a word, the first or last known instance/example of an object/phenomenon/etc., or a simple list of examples or facts - are not allowed as standalone threads.
Depth
@@
An in-depth answer provides the necessary context and complexity that the given topic calls for, going beyond a simple cursory overview. Your answer should be giving context to the events being discussed, not simply listing some related facts.
Sources
@@
We do not require sources to be preemptively listed in an answer here, but do expect that respondents be familiar with relevant and reliable literature on the topic, and that answers reflect current academic understanding or debates on the subject at hand. But sole reliance on tertiary sources for context and analysis is not allowed, and will result in the removal of a response.
No personal anecdotes
Personal anecdotes are not acceptable answers in this subreddit.
No speculation
@@
Suppositions and personal opinions are not a suitable basis for an answer here. Warning phrases for speculation include: "I guess…" or "My guess is…", "I believe…", "I think…", "… to my understanding", "It makes sense to me that…", "It’s only common sense."
No partial answers or "placeholders"
@@
An answer should be full and complete in and of itself.
No political agendas or moralizing
@@
This subreddit is a place for learning and open-minded discussion. As such, answers should not be written in the interests of advancing a personal agenda, but should represent a sincere effort to make an argument from the historical record. They should be constructed in keeping with the principles of the historical method - that is to say, your evidence should not be chosen selectively to support an argument that you feel is right; your argument should instead demonstrably flow from your critical engagement with an appropriate range of evidence.
Do not just post links or quotations
@@
Do not just post links to other sites as an answer. This is not helpful. The expectation is that a user is posting to this subreddit because they are looking for the type of answer dictated by the rules in place here. Please take some time to put the links in context for the person asking the question. Avoid only recommending a source – whether that’s another site, a book, or large slabs of copy-pasted text. If you want to recommend a source, please provide at least a small summary of what the source says.
No plagiarism
@@
We have a zero-tolerance policy on blatant plagiarism, such as directly copying and pasting another person’s words and trying to pass them off as your own.
Jokes and humor
@@
A post should not consist only of a joke, a humorous remark, or a flippant comment. You can certainly include humor as part of a full and comprehensive post, but your post should not be made solely for the purpose of being funny.
Digression
@@
All comments should be related to the topic as outlined in the original post.
Clutter
@@
Please refrain from commenting for the sake of commenting. This includes, but is not limited to, statements about how interesting the question is, how you would like to see an answer, to "remember to come back later," to share a story that the question reminds you of, and so on.
@@
No Trolling
trolling; provocative
TRAC2020 _IBEN_A_ bert-base-multilingual-uncased
No
Yes
Yes
Personal Attacks or Flaming
doxxing; cyberbullying
TRAC2020 _IBEN_A_ bert-base-multilingual-uncased
No
Yes
Yes
Off Topic
topic classifier
tweet-topic-21-multi
No
No
Yes, but with changes
Twitter topic classification
Need to be finetuned on this topic
@@
Spam
spam
bert-tiny-finetuned-sms-spam-detection
No
No
Yes, but with changes
Detect message spam
1. Need to be adapted to reddit texts rather than message texts. 2. Need to be adapted to detect self-promotion of some youtube account/website etc.
Low-Effort Posts
low effort; low quality
NA
No
No
No
Proselytizing
proselytizing; Soapbox
NA
No
No
No
Harassment or Bigotry
harassment; bigotry
Text-Moderation
No
Yes
Yes, but with changes
any and all curse words are permitted
Brigading
brigading
NA - Toxicity
No
Yes
Yes, but with changes
These models may be able to detect some of the brigading comments, but there is no specific model for brigading and some brigading encouraging comments may not be caught by the toxicity models
@@
Incivility - No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhistling/ Homophobic Slurs
offensive; hatespeech
Text-Moderation
No
Yes
Yes, but with changes
Yes
Quoting something offensive from a movie is okay but needs to be put in quotation marks.
No Spam & Self-promotion - Same Source Posts
spam
bert-tiny-finetuned-sms-spam-detection
No
No
Yes, but with changes
No
Detect message spam - 1. Need to be adapted to Reddit texts rather than message texts. 2. Need to be adapted to detect self-promotion of some YouTube account/website etc.

### v3

@@ 1 Introduction
In this paper, we aim to assess to what extent current natural language processing (NLP) models can serve the wide spectrum of moderation rules so that they can be helpful in assisting moderators. First, to understand the functions previous automated content moderation models have focused on, we conduct a model review on Hugging Face (HF) with rules from three subreddits as exemplars. This allows us to gauge the progress of past model developments in covering various moderation rules. We use model review as opposed to the more common literature review to gain a technical understanding of the existing models’ functions. In addition to examining models that are built to handle specific tasks, we also assess so-called “general-purposed” large language models’ (LLMs’) capability in covering various moderation rules. We evaluate GPT-4 and Llama-2 on a new evaluation dataset that we collected, consisting of moderation decisions from r/AskHistorians and covering a wide range of rules. Finally, using the models’ performance on this new dataset as an empirical grounding, we conducted a survey study with active moderators from r/AskHistorians ($N=11$). Through this survey study, we aim to gain insights into model users’ perspectives on the performance requirements for useful moderation models on different rules.
We find a substantial gap in both existing NLP models and LLMs’ performance when to cover the diverse set of moderation rules that subreddits employ. Our analysis shows the majority of moderation rules from the three subreddits ($\sim 80\%$) are unrelated to toxicity detection, with nearly $70\%$ lacking an huggingface model designed for their resolution. While one might hope that general-purpose LLMs could fill this gap, our experiments with GPT-4 and Llama-2 show that LLMs fall short in their ability to detect violations of many rules. Specifically, both GPT-4 and Llama-2 exhibited moderate to low precision and/or recall ($<70\%$) for half of the rules from r/AskHistorians. Findings from our survey study also indicate that neither LLM has good enough performance to be useful for $6$ of $23$ r/AskHistorians rules ($26\%$), including rules such as Scope, Digression, and Sources [^3] For explanation of the rules, please refer to Table 2 in appendix.. Meanwhile, our survey study also shows that moderators are excited about an assistant model—they are okay with even imperfect tools if they are well-informed about their limitations. For different kinds of rules, moderators have complex needs in terms of model precision and recall. For instance, they need high recall for complex rules (e.g. Plagiarism and Digression) and high precision for simple rules (e.g. Current Event and Jokes and Humour). Our study highlights the necessity of, and provides a guide for, future content moderation work to expand its focus beyond toxicity detection, encompassing a broader spectrum of moderation rules to meet the needs of moderators seeking automation assistance.
@@ 2 Background and Related Work
Recently, Kumar et al. (2023) tested LLMs’ performance on rule-based moderation with rules from $95$ subreddits. They evaluated GPT models (3, 3.5, 4) on a dataset consisting of removed Reddit posts as violating and not violating content. They conclude that the GPT models are effective in conducting moderation on subreddits like r/Movies but struggle with other subreddits such as r/askscience and r/AskHistorians. Our study builds on previous results by studying gaps in both purpose-built moderation models and LLMs. Rather than focusing on the broad binary questions of removing or keeping posts, we focus on coverage for specific moderation rules. In addition, we include the perspectives of human moderators to understand how existing models address the needs of real-world moderation.
@@ 3.1 Method
We then search on HF by matching keywords with model names or model cards to find the most relevant model to each moderation rule. We traverse the top 20 search results to find the most relevant model, which we call the matching model. If more than one model is relevant, we select the model with the highest number of likes on HF. If nothing relevant is found, we declare there is no matching model for this moderation rule. We also skip any model that does not contain a model card (i.e., no model description). See Tables 3, LABEL:tab:annotation_movies, and LABEL:tab:annotation_ask in the appendix for the keywords we used and the matching models for each rule.
@@ 3.2 Model Review Results
Among the $37$ non-toxicity-related rules, only $6$ ($20\%$) rules have matching HF models that can be applied directly. To see some evaluation results of these matching models, please refer to Appendix A in appendix. $13$ ($43\%$) of the rules require some model modifications in order for the matching model to be adapted for the rule, whereas $11$ ($37\%$) rules have no matching model at all. These rules include, for example, low-effort posts and proselytizing from r/Atheism, and no homework and no ‘‘soapboxing’’ or loaded questions from r/AskHistorians.
@@
To see detailed annotations for each question, please refer to Tables 3, LABEL:tab:annotation_movies, and LABEL:tab:annotation_ask in appendix.
@@ 4.1 Method
The evaluation dataset was created and annotated by a volunteer moderator for r/AskHistorians who has expertise in both moderation and qualitative analysis. For each rule, 1111111113131313 violating posts or comments were selected, save for posts violating the Illegal artifacts rule, which, due to the rareness of violations, was only represented $3$ times in the dataset. Most content was selected from r/AskHistory, a Reddit community that is thematically similar to r/AskHistorians but with different moderation rules. Some additional content was taken directly from r/AskHistorians. We exclude all borderline cases and only select posts and comments that clearly violate the rules. In total, the dataset includes $101$ questions and $134$ comments. [^5] In r/AskHistorians, posts are questions users have asked and comments are responses to questions. Comments included in the dataset are not necessarily responses to the questions included in the dataset. Because content often violates more than one rule, questions and comments included in the dataset often reflect multiple violations; however, annotations represent the most relevant or severe violation. For example, a comment containing only a joke would be annotated as violating the rule prohibiting joke responses, even though it would also violate rules on comprehensiveness or depth. The dataset also included $11$ questions and $10$ comments that do not violate any r/AskHistorians rules. We use these data as our evaluation dataset. [^6] The dataset, along with a datasheet, are available under a MIT licence at: https://github.com/TristaCao/into_inclusivecoref.
@@
Finally, following the suggestions from Shieh (2023), we set the model’s role as a moderator assistant and put the main question at the beginning of the prompt. In the few-shot setting, we also repeat the main question at the end of the prompt. See Appendix B for the exact prompts we used.
@@ 4.2 Moderator Evaluation
To understand volunteer moderators’ perspectives on the relative performance of LLMs, we also conducted an evaluation survey with active moderators from r/AskHistorians [^7] Approved by our institutional IRB, #1704882-9.. The survey aimed to capture 1) whether moderators would use the LLM to help with their moderation work on each rule, given the LLM’s performance (precision and recall) identifying violations for this rule, and 2) for each rule, what kind of model performance is needed to effectively assist moderators.
@@ 4.3 Model Performance – Results
Our survey results indicate that different moderators have different perceptions of what constitutes an adequate model: for the evaluation questions, participants exhibited only fair agreement (Fleiss’ Kappa $0.34$). Mirroring policy decision-making standards in the subreddit, we use majority vote to summarize participants’ evaluation of the models’ performance. The results are shown in Figure 2 and Figure 3. In most cases, moderators would consider models with $>70\%$ precision and recall to be useful. Among the $23$ rules, both models are considered useful for three question rules and three comment rules. For some rules, such as NSFW/Sensitive content, No plagiarism, No political agendas or moralising, and Depth, moderators can accept low recall; for some, such as Do not just post links or quotations, moderators can accept low precision. Importantly, there are six rules ($26\%$ of all rules) for which neither model is considered useful. These rules are Homework, Privacy, and Scope question rules and No partial answer or ‘‘placeholders’’, Digression, and Sources comment rules. This emphasizes that, while current LLMs provide possibilities for supporting content moderation work, further improvement is needed to support useful moderation tools.
@@ 5 Discussion
Furthermore, we observed varying requirements among users and rules. Even for the same rule, different moderators expressed distinct preferences for high precision or recall. This suggests that in situations where a trade-off between precision and recall is necessary, moderators may lean towards a model offering more user control, thus allowing for an adjustable trade-off. This way, they can customize the model to align with their preferences. In addition, our participants also expressed some common preferences. For complex rules (e.g. Plagiarism and Digression), moderators prefer higher recall and model explanations over mere flagging of violations. Conversely, for simpler rules (e.g. Current Event and Jokes and Humour), they prefer higher precision.
@@ 7 Limitations
Appendix A Evaluation of HF Matching Models
We evaluated the HuggingFace matching models with our r/AskHistorians evaluation dataset. As indicated in LABEL:tab:annotation_ask, there are four non-toxicity-related rules that have matching HF models: NSFW/Sensitive content, Illegal artifacts, Basic Facts, and Jokes and humor. Among these four matching models, the model for Illegal artifacts is not compatible with the HuggingFace model pipeline and lacks the appropriate code for inference. Similarly, the model for Basic Facts uses a model structure that is not defined within Huggingface module.
Consequently, we evaluated only the models for NSFW/Sensitive content and the model for Jokes and humor using our dataset. The performance of the NSFW/Sensitive content model is: precision $57.1\%$ and recall $92.3\%$. The performance of the Jokes and humor model is: precision $55.6\%$ and recall $45.5\%$.
Appendix B Prompt for Testing LLMs
B.1 Prompt for Testing Llama-2 with the Few-shot Setting
@@
B.2 Prompt for Testing GPT-44 with the Zero-shot Setting
@@
Appendix C LLMs Performances under Zero-shot and Few-shot Settings
@@
Appendix D User Survey
@@
Advertising & Commercialization No Spam & Self-promotion - Same Source Posts
@@
Content/Behavior Homework; No "what if" questions Basic facts Depth; No speculation; Jokes and humor Proselytizing No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs; No Extraneous Comic Book Movie Submission"
Copyright No plagiarism
Doxxing/Personal Info Privacy; No personal anecdotes Personal Attacks or Flaming
Format No partial answers or "placeholders"; Do not just post links or quotations No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title
Harassment Civility Harassment or Bigotry No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
Hate Speech Civility No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
@@
Links & Outside Content Sources Prohibited Popular Websites
Low-Quality Content Scope; Clutter Low-Effort Posts
NSFW NSFW/Sensitive content
Off-topic Digression Off Topic No Extraneous Comic Book Movie Submission
Personal Army Brigading No Subreddit Brigading
Personality Civility Personal Attacks or Flaming No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhisteling/ Homophobic Slurs
Politics No political agendas or moralising; No "Soapboxing" or loaded questions
@@
Reposting No Repost or Discussion Threads of New Releases
@@
Spam Spam No Spam & Self-promotion - Same Source Posts
Spoilers No Repost or Discussion Threads of New Releases; No Spoilers
Trolling No Trolling No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title
Voting "Poll-type" question
Others Current event; Illegal artifacts Don’t complain about the use of AAVE or slang No Encouraging Piracy
@@
Civility All users are expected to behave with courtesy and politeness at all times. We will not tolerate racism, sexism, or any other forms of bigotry. This includes Holocaust denialism. Nor will we accept personal insults of any kind, and do not allow minor nitpicking of grammar or spelling.
@@
Scope Submissions must be about a question about the human past, a META post about the state of the subreddit, or an AMA ("Ask Me Anything") with a historical expert or panel of experts.
@@
No Current events To discourage off-topic discussions of current events, questions, answers, and all other comments must be confined to events that happened 20 years ago or more, inclusively (e.g., 2003 and older).
@@
Homework Our users aren’t here to do your homework for you, but they might be willing to help. Don’t just give us your essay/assignment topic and ask us for ideas.
@@
NSFW/Sensitive content Questions with NSFW titles will be deleted, and we will ask you to repost it with a different title.
@@
No "Poll"-type questions "Poll"-type questions aren’t appropriate here: "Who was the most influential person in history?" or "Who was the worst general in your period?" or "Who are your Top 10 favorite people in history?" If your question includes the words "most" or "least", or "best" or "worst" (or can be reworded to include these words), it’s probably a "poll"-type question.
@@
No "Soapboxing" or Loaded Questions All questions must allow a back-and-forth dialogue based on the desire to gain further information and not be predicated on a false and loaded premise in order to push an agenda. Example: Good Question: "People say that Nixon is the worst President of all time. Why is this so?" Bad Question: "Nixon was the worst President of all time. Why isn’t Obama considered the worst?" The bad question is a fishing expedition to try to start a debate about Obama’s presidency. Most of these questions will break our 20-year rule, or try to set up a debate about an issue using a long wall of text in the main post.
@@
No "What If" Questions Questions should be about what did happen, not what could have happened.
@@
Privacy We do not allow questions that pose possible privacy issues for living or recently deceased persons who are not in the public eye. The cut-off for "recent" is 100 years.
@@
Illegal artifacts It is our policy to disallow posts asking for further information on artifacts where there is a likelihood that the acquisition or possession of the item might be illegal, unethical, and/or run contrary to sound, historical practices.
@@
Basic Facts Questions looking for specific, basic facts - for the purpose of this rule, seeking a name, a date or time, a number, a location, the origin of a word, the first or last known instance/example of an object/phenomenon/etc., or a simple list of examples or facts - are not allowed as standalone threads.
@@
Depth An in-depth answer provides the necessary context and complexity that the given topic calls for, going beyond a simple cursory overview. Your answer should be giving context to the events being discussed, not simply listing some related facts.
@@
Sources We do not require sources to be preemptively listed in an answer here, but do expect that respondents be familiar with relevant and reliable literature on the topic, and that answers reflect current academic understanding or debates on the subject at hand. But sole reliance on tertiary sources for context and analysis is not allowed, and will result in the removal of a response.
No personal anecdotes Personal anecdotes are not acceptable answers in this subreddit.
@@
No speculation Suppositions and personal opinions are not a suitable basis for an answer here. Warning phrases for speculation include: "I guess…" or "My guess is…", "I believe…", "I think…", "… to my understanding", "It makes sense to me that…", "It’s only common sense."
@@
No partial answers or "placeholders" An answer should be full and complete in and of itself.
@@
No political agendas or moralizing This subreddit is a place for learning and open-minded discussion. As such, answers should not be written in the interests of advancing a personal agenda, but should represent a sincere effort to make an argument from the historical record. They should be constructed in keeping with the principles of the historical method - that is to say, your evidence should not be chosen selectively to support an argument that you feel is right; your argument should instead demonstrably flow from your critical engagement with an appropriate range of evidence.
@@
Do not just post links or quotations Do not just post links to other sites as an answer. This is not helpful. The expectation is that a user is posting to this subreddit because they are looking for the type of answer dictated by the rules in place here. Please take some time to put the links in context for the person asking the question. Avoid only recommending a source – whether that’s another site, a book, or large slabs of copy-pasted text. If you want to recommend a source, please provide at least a small summary of what the source says.
@@
No plagiarism We have a zero-tolerance policy on blatant plagiarism, such as directly copying and pasting another person’s words and trying to pass them off as your own.
@@
Jokes and humor A post should not consist only of a joke, a humorous remark, or a flippant comment. You can certainly include humor as part of a full and comprehensive post, but your post should not be made solely for the purpose of being funny.
@@
Digression All comments should be related to the topic as outlined in the original post.
@@
Clutter Please refrain from commenting for the sake of commenting. This includes, but is not limited to, statements about how interesting the question is, how you would like to see an answer, to "remember to come back later," to share a story that the question reminds you of, and so on.
@@
No Trolling trolling; provocative TRAC2020 _IBEN_A_ bert-base-multilingual-uncased No Yes Yes
Personal Attacks or Flaming doxxing; cyberbullying TRAC2020 _IBEN_A_ bert-base-multilingual-uncased No Yes Yes
Off Topic topic classifier tweet-topic-21-multi No No Yes, but with changes Twitter topic classification Need to be finetuned on this topic
@@
Spam spam bert-tiny-finetuned-sms-spam-detection No No Yes, but with changes Detect message spam 1. Need to be adapted to reddit texts rather than message texts. 2. Need to be adapted to detect self-promotion of some youtube account/website etc.
Low-Effort Posts low effort; low quality NA No No No
Proselytizing proselytizing; Soapbox NA No No No
Harassment or Bigotry harassment; bigotry Text-Moderation No Yes Yes, but with changes any and all curse words are permitted
Brigading brigading NA - Toxicity No Yes Yes, but with changes These models may be able to detect some of the brigading comments, but there is no specific model for brigading and some brigading encouraging comments may not be caught by the toxicity models
@@
Incivility - No Antagonism/ Flamewars/ Attention Whoring - No Racial/ Sexist/ Dogwhistling/ Homophobic Slurs offensive; hatespeech Text-Moderation No Yes Yes, but with changes Yes Quoting something offensive from a movie is okay but needs to be put in quotation marks.
No Spam & Self-promotion - Same Source Posts spam bert-tiny-finetuned-sms-spam-detection No No Yes, but with changes No Detect message spam - 1. Need to be adapted to Reddit texts rather than message texts. 2. Need to be adapted to detect self-promotion of some YouTube account/website etc.